you have $500 dollars in saving account. One year from now, you have $625 in that savings account. Write an equation that models the situation. How many years until you have $1,000.
Answer by addingup(3677)

500(1+x)^1 = 625
500 + 500x = 625
500x = 125
x = 0.25
interest = 0.25 * 100 = 25%
----------------------------
How many years until you have 1000:
- what is the interest rate?
- Is the interest compound or simple?